Circular No. 8608

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 S3 (READ) Michael T. Read reports his discovery of a comet (tabulated below) on

CCD images taken with the 0.9-m f/3 Spacewatch reflector, the object being diffuse with a long tail toward the southwest. Several other observers have noted the object’s cometary appearance on their CCD images, including J. Lacruz (La Cañada, Spain, 0.30-m reflector, Oct. 1.0 UT; 22′′ tail in p.a. 250^{o}), P. Birtwhistle (Great Shefford, U.K., 0.40-m reflector, Oct. 1.11– 1.16; very concentrated nuclear condensation 4′′ in dia, with a faint coma of diameter 8′′ and a straight, thin tail 48′′ long in p.a. 252^{o}), E. J. Christensen (Mt. Lemmon, 1.5-m reflector, four 30-s co-added exposures, Oct. 1.4; wellcondensed 5′′ coma of mag 18.5 with a 25′′ tail in p.a. 240^{o}–250^{o}), J. E. McGaha (Tucson, 0.62-m reflector, Oct. 1.4; 4′′ × 6′′ coma elongated with narrow, curved 20′′ tail in p.a. 245^{o}), and J. Young (Table Mountain, 0.6-m reflector, Oct. 1.4; bright diffuse coma of diameter 6′′ and 20′′ fan-shaped tail spanning p.a. 245^{o}–265^{o}).

2005 UT  α_{2000}  δ_{2000}  Mag.
Sept.30.41409  2^{h}27^{m}29.^{s}77  +19^{o}18′27′′.3  19.2

The available astrometry and the following preliminary parabolic orbital elements appear on MPEC 2005-T11; the comet is likely of short period.

T = 2005 May 16.916 TT  ω = 55.^{o}964
Ω = 288.535 }2000.0
q = 2.42927 AU  i = 3.139

SUPERNOVAE 2005eq, 2005er, AND 2005es Further to IAUC 8605, N. Ponticello, E. Lee, and W. Li report the

LOSS discovery of three apparent supernovae on unfiltered KAIT images:

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005eq Sept. 30.40 3^{h}08^{m}49.^{s}31 − 7^{o}01′59′′.7 17.4 15′′.9 E, 26′′.3 N
2005er Sept. 29.22 22 50 00.84 +11 37 05.7 18.7 91′′.1 E, 34′′ N
2005es Sept. 30.25 23 25 47.62 + 8 02 17.3 18.4 2′′.6 W, 8′′.3 S

Additional magnitudes: SN 2005eq in MCG –01-9-6, Sept. 15.39 UT, [19.0; Oct. 1.39, 17.2. SN 2005er in NGC 7385, Sept. 8.21, [19.5; 13.23, [19.0; 24.22, [18.0; Oct. 1.23, 18.4. SN 2005es in MCG +01-59-79, Sept. 11.23, [19.5; 21.24, [18.0; Oct. 1.26, 18.7.

2005 October 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green